# Aptly: Making Mobile Apps from Natural Language

### v1

@@
Abstract
We present Aptly, an extension of the MIT App Inventor platform enabling mobile app development via natural language powered by code-generating large language models (LLMs). Aptly complements App Inventor’s block language with a text language designed to allow visual code generation via text-based LLMs. We detail the technical aspects of how the Aptly server integrates LLMs with a realtime collaboration function to facilitate the automated creation and editing of mobile apps given user instructions. The paper concludes with insights from a study of a pilot implementation involving high school students, which examines Aptly’s practicality and user experience. The findings underscore Aptly’s potential as a tool that democratizes app development and fosters technological creativity.
Index Terms:
@@ Abstract
MIT App Inventor, Large Language Model, Block programming
§§footnotetext: Equal contribution
I Introduction
@@
The rise of technology in daily life highlights the need for everyone, particularly youth, to engage with digital tools to solve real-world challenges. Mobile applications serve as a powerful medium for such engagement given the widespread adoption of smartphones and tablets across diverse demographics, including under-represented communities [1]. However, the traditional app development process, which often requires a solid foundation in computer science, has historically excluded many aspiring creators from participating. To democratize this process, the MIT App Inventor initiative has revolutionized the creation of mobile apps by introducing a user-friendly, visual programming environment that empowers anyone, regardless of their technical background, to develop functional applications [2]. App Inventor utilizes a block-based programming approach, which allows users to visually assemble their applications using geometric shapes, making the development process more intuitive and accessible [3]. The tool not only fosters creativity among novice programmers but also encourages a more inclusive approach to technological education and development [4]. Tissenbaum et al. [5] call the process by which people leverage technology like App Inventor to make societal impact computational action.
@@
The advent of large language models (LLMs) and their ability to generate code [6, 7] has opened new avenues for mobile app development. Capitalizing on this technological advancement, our research introduces Aptly, a groundbreaking platform that allows users to create mobile applications using natural language to lower barriers to computational action. For example, a user could simply state, “Make me an app that translates English to Spanish,” and Aptly would automatically generate a functional app based on this input. This paper details the framework of Aptly, beginning with an overarching view of the system’s architecture. We will discuss the development of the Aptly Language and our method of integrating it with an LLM to produce the desired syntax. Additionally, we explore the mechanisms that enable users to refine their apps through natural language commands. We conclude by examining Aptly’s current functionalities and presenting preliminary findings from its deployment among high school students.
II Aptly Architecture
Figure 1: An interaction diagram between the different components of the MIT App Inventor with Aptly system.
Aptly augments MIT App Inventor with generative AI coding capabilities. We accomplish this by building on a version of App Inventor supporting real-time collaboration (RTC) [8]. The Aptly agent acts as a collaborator with users, taking their instructions in natural language and manipulating the project just as another user might. The overall system involves the following components:
  • MIT App Inventor: The base version of the App Inventor [^1] https://github.com/mit-cml/appinventor-sources is extended to integrate with the Aptly system and the RTC server to facilitate enhanced functionality.
  • MIT App Inventor RTC: The real-time collaboration (RTC) server mediates between agents in the system architecture.
  • Aptly Server: Detailed in the subsequent section, the Aptly server invokes the LLM and converts its responses into events understood by App Inventor RTC.
  • OpenAI GPT-3.5/4: Utilizing OpenAI’s LLM models, this component provides advanced code completion services [9, 10].
Figure 1 depicts the relationship between the various components as they are used by the app creator.
II-A The Aptly Language
@@
Figure 2: An example of the Aptly code in Listing 1 and its correspondence to the App Inventor interface.
@@
Given these attributes, we hypothesize that generative models can be effectively fine-tuned to interface with Aptly, leveraging its Python-like structure. However, it is important to note that while Aptly draws inspiration from Python, it is distinct. Aptly has been specifically constrained to ensure a one-to-one correspondence with App Inventor: every valid App Inventor program is also a valid Aptly program, and vice versa. See Listing 1 for an example Aptly program and Figure 2 for showcasing how the code is represented as blocks.
Figure 3: When the user requests an application with its textual description, we automatically synthesize the prompt by adding several example pairs along with the desired application’s textual description. This constructed prompt is fed into one of OpenAI’s GPT-X models as an input, which outputs code that can be converted into a fully functional mobile application
To ensure the fidelity and accuracy of the one-to-one correspondence between the Aptly and App Inventor languages, we employ a comprehensive suite of integration tests. These tests validate the conversion cycle by processing Aptly code through a round-trip transformation to an App Inventor (.aia) file and back to Aptly. Furthermore, any new additions to the Aptly language introduced through pull requests are required to include corresponding tests that affirm the functionality and behavior of the new features.
For maintaining the precise correspondence of the Aptly language, we define the following four components:
  • Parser: Parses the Aptly code into its respective representation in the Abstract Syntax Tree (AST). Each method adheres to the extended Backus-Naur Form (eBNF) of the Aptly language.
  • ProjectReader: Transforms the App Inventor representations into the corresponding Aptly AST.
  • ProjectWriter: Converts Aptly AST into the App Inventor representation (either in the designer or blocks format).
  • Serializer: Outputs the AST into the Aptly code format.
In tandem, these four components provide the core functionality to enable app creation with Aptly by allowing us to convert between the text and block representations via the Aptly AST.
Figure 4: An app created from a prompt given in Dutch.
II-B Few-shot Prompt Engineering
Aptly leverages machine learning research on generative models, which are probabilistic frameworks that learn how datasets are generated to facilitate new data creation. We specifically employ a transformer-based large language model to develop mobile apps [11]. This approach capitalizes on the recent significant advancements in large language models, an area that has garnered substantial interest for its impressive research outcomes and substantial commercial opportunities. Generative models operate based on “prompts” which are crucial inputs that significantly influence their performance. The effectiveness of these models hinges on the quality of the prompts, a practice known as prompt engineering—an active field within machine learning research [12]. A common method of prompt engineering involves “few-shot” prompts, where a model is given a limited number of solved task examples as part of its input. This approach is often interpreted as the model “learning” the task in real-time from these few examples [9]. Aptly uses Open AI’s LLMs, which are capable of generating computer code from natural language descriptions in several programming languages including Java, JavaScript, Python, C, C#, and Swift.
@@ II-B Few-shot Prompt Engineering
Aptly’s method for providing relevant prompts to accompany a description $D$ is to provide a set of example application descriptions paired with appropriate Aptly Code. We use semantic similarity to highlight the “appropriateness” of code and description. The method relies on having a way to automatically measure the similarity of text descriptions and the program. For computing the similarity of the text description and program, we use embeddings of data elements, which are numerical representations of concepts converted to number sequences [13]. In our scenario, an embedding represents the semantic meaning of a natural language description or code. To measure the similarity of two items, one embeds them in the same vector space and takes the distance between the two embedding vectors as a measure of similarity (Small distance implies high similarity while large distance implies low similarity). We use cosine distance, which reflects the angle between vectors, to compute the similarity between two vectors.
@@
We employ OpenAI’s Text/Code Embedding model to generate embeddings from descriptions and associated code. These neural network models utilize Contrastive Pre-Training—a training method that functions by clustering in vector space [14]. Specifically, it draws predefined positive examples (i.e., matching text and code) closer together, while pushing negative examples (i.e., contrasting text and code) apart. This technique enhances our model’s ability to identify and select the example pair most relevant to a user’s specified description. We feed the synthesized prompt into GPT, which outputs the Aptly Code corresponding to the user’s requested app description. We can then convert the generated code into App Inventor blocks to generate a fully functional application (Figure 3).
An interesting feature of the use of LLMs for this process is that the user input need not be expressed in English as might be the case with traditional natural language processing techniques. As an example, we provided this prompt in Dutch: “Maak een app met een tekstbox, een lijst met zes talen en een knop met de tekst ‘vertalen’. Wanneer er op de knop wordt gedrukt dan vertaal je de tekst in de geselecteerde taal en laat de vertaling zien.” This produced an app that gave the user the option to translate into English, Spanish, French, German, Italian, and Japanese, with all of the app content expressed in Dutch (Figure 4). Importantly, all of the example pairs we provide the LLM are written in English–there are no examples expressed in Dutch. Ultimately, these abilities of LLMs will reduce the barrier to computational action by removing the need to learn English before or while learning to code.
Figure 5: An overview of the information flow during an edit operation using Aptly. Arrows represent the flow of information and should be read as “tail label head,” e.g., “App Inventor sends AIA file” or “RTC Event Sequence consumed by App Inventor RTC.”
II-C The Editing Capability
Often it is unreasonable to assume one can create a full app in a single shot request. We aim to empower programmers not only to generate apps using natural language, but also to refine, iterate, and evolve their designs continuously through natural language commands. This approach allows for a more dynamic development process, enabling creators to adapt and modify their applications easily as new ideas or requirements emerge. This section details how editing of apps is achieved in Aptly.
Figure 6: The update operations (blue) to transform the AST of the original program into the AST of the modified program. Some unmodified nodes are omitted for brevity.
II-C1 Editing Projects using Descriptions
Editing a project with Aptly is done by emitting a sequence of edit operations over the MIT App Inventor RTC system. When a user requests an edit by providing a description of their desired change, App Inventor sends the AIA for the current project plus the user-provided edit description to the endpoint of the Aptly server.
The editing process uses the following steps (see also Figure 5). The Aptly frontend sends the project state and user request to the server (1), which utilizes the ProjectReader and Serializer to obtain the corresponding Aptly code. The server then constructs an edit prompt–a top-ranked selection of example edits plus the project (converted to Aptly code) and the description–to be sent (2) to the LLM. We use the same technique for ranking examples as discussed previously.
When the LLM returns updated Aptly code, the Aptly server constructs the AST representation of the original code (via the ProjectReader) and the modified code (via the Parser), and computes the minimum-cost sequence of edit operations to transform the original AST into the modified one. The Aptly server utilizes the tree edit distance algorithm proposed by Zhang and Shasha [15] to compute the minimum-cost sequence of edit operations (Fig. 5 note 3). Aptly specifically utilizes the Python zss module.The edit operations returned are INSERT, REMOVE, UPDATE, and MATCH. The update_cost function is customized for App Inventor by taking into account the type of two nodes; for example, if there are two component declarations (e.g., Button1 = Button(...) and Label1 = Label(...)), there should be infinite update cost to transform one of those component declarations into the other because Button and Label are two different types and App Inventor does not allow a way to change a button into a label or vise versa. Similarly, in App Inventor, if there are two bound component events (e.g., when Button1.Click() and when Button2.Click()), it should only be possible to update one to be the other if the type of the component is the same because this is the only type of change App Inventor allows in the Blocks editor for a bound component event. If the user wants to change which method they’re using (e.g., Click vs. TouchUp) or change the event to a different component type (e.g., Button1.GotFocus vs Switch1.GotFocus), the user would have to grab a new block from the blocks drawer and delete the old block, so Aptly cannot make this change via an UPDATE operation either.
Given the sequence of insertions, removals, updates, and matches recommended by zss, the TreeDiff class emits a sequence of edit events compatible with the App Inventor RTC system (Fig. 5 note 4). These edit events include creating, updating, moving, and deleting components and blocks. Some changes require multiple edit events; for example, since block stacks are represented as linked lists, inserting a block in the middle of a block stack requires creating the block, moving it to the correct location (i.e., setting its parent correctly), and moving its descendent (if applicable) to now point to it rather than its previous parent. As another example, deleting an item in a list or deleting a pair in a dictionary requires issuing a move event for all items after the deleted item in the list, as even though these structures are not represented as linked lists, each subsequent item’s input name (e.g., ‘ITEM1’ vs. ‘ITEM0’ for list items) will change.
To emit the correct edit events via the RTC system, we process the operations returned from zss in several stages:
  • Preinsertion: To handle moving components, we need to perform a preprocessing step that aggregates all of the new components into a map based on their name and type. We process all of the insert operations to build this map, which will be consumed as part of the remove stage to determine whether a component was renamed, moved, or removed. If we did not include such a stage, we might end up deleting and then inserting the same node rather than moving a node, incurring unnecessary cost.
  • Remove: We emit removal events for deleted components first. When a component is removed, we log this information so that any block removals involving that component are ignored since they are covered by the component’s removal. Likewise, we need to log all of the block IDs that are removed so that we do not also emit removal events for the subtree starting at that node. Note that there is a special case in regards to moved components; for subtrees starting at nodes that are marked as insertions, a removal does not take place.
  • Match: The Match operation in the tree edit sequence is a no-op. However, when translating the operations into App Inventor events, we use this step to perform some bookkeeping necessary to support insertion operations later.
  • Update: The Update stage converts the value of a node in the source tree to a new value at the same position in the destination tree. Common reasons for update operations involve renaming components, changing property values, or renaming variables or procedures.
  • Insert Stage: The Insert operation adds a new subtree to the AST. Because we enumerate the changes in breadth-first order, an insertion in the tree covers the whole subtree at that node. When subtrees are inserted, we must return to looking at the parent node (which would be a match/update operation) in order to issue the necessary move events needed based on the inserted node (such as moving the new node to the right place and adjusting the parent and input name of successors if needed).
Once all of the edit events have been created, the Aptly server joins the user’s project in the RTC system and emits the stream of changes (Fig. 5 notes 5, 6).
II-C2 Applying Updates via Real-Time Collaboration System
Aptly leverages the App Inventor Real-Time Collaboration (RTC) [8] system to edit the project based on the event sequence produced in the previous step. Aptly is conceptualized as a collaborative user within the project environment. Once the event sequence is generated by the Aptly server, it joins the user’s project in the RTC server. It then sends the event sequence to the server, which relays the changes to the user’s App Inventor session to be applied to the project. Leveraging the RTC system, Aptly allows for modifications to be reversed if they do not meet the user’s expectations. This functionality is facilitated by the RTC’s capability to execute events in reverse, relying solely on a stack of events that can be undone, thereby eliminating the need for maintaining additional state information. This design enhances user control and flexibility, making the editing process more user-friendly and adaptable to changes.
II-C3 Worked Example
Consider the following Aptly code representing the app “HelloPurr,” App Inventor’s equivalent to a Hello World program.
Screen1 = Screen(AppName = "HelloPurr")
Cat = Button(Screen1, Image = "kitty.png")
Meow = Sound(Screen1, Source = "meow.mp3")
when Cat.Click():
call Meow.Play()
One might offer in natural language “add a label below the kitty that says ‘pet the kitty’.” The LLM may generate the following Aptly code in response:
Screen1 = Screen(AppName = "HelloPurr")
Cat = Button(Screen1, Image = "kitty.png")
Label1 = Label(Screen1, Text = "pet the kitty")
Meow = Sound(Screen1, Source = "meow.mp3")
when Cat.Click():
call Meow.Play()
An illustration of the above two codes AST, respectively noted as $a$ and $b$, is shown in figure 6. The two ASTs are then processed into a sequence of edits needed to transform the AST of the first snippet into the AST of the second snippet. In this particular scenario, the Zhang-Shasha algorithm will result in MATCH operations except for the INSERT operations that introduce the AST nodes corresponding to the insertion on line 3:
INSERT Label1 (Identifier)
INSERT Label (Reference)
INSERT Screen1 (Reference)
INSERT Text (Reference)
INSERT "pet the kitty" (StringValue)
INSERT Text = "pet the kitty" (DesignerAssignment)
INSERT ComponentTypeInvocation(Label1, Screen1, {Text: "pet the kitty"}) (ComponentTypeInvocation)
INSERT ComponentDecl(Label1, ComponentTypeInvocation(Label1, Screen1, {Text: "pet the kitty"})) (ComponentDecl)
For the processing of INSERT operations, Aptly considers the operations in a breadth-first fashion, which leads to the ComponentDecl being the first node, n, encountered during traversal. We know that an insertion of a node a depth d subsumes all of the children below depth d in the subtree rooted at n. The TreeDiff class checks to see if it has a method to handle the insertion of the AST type of n (i.e., ComponentDecl). If such a method exists, it is called. Otherwise, the subtree starting at n will continue to be traversed breadth-first. In the case of ComponentDecl, such a method does exist and it interprets the insertion of a ComponentDecl node as the RTC event sequence of ComponentAdd, ComponentMove, and one or more ComponentProperty events. For example, if the UUID of the label component is 12345:
ComponentAdd(12345, "Label", [])
ComponentMove(12345, 0, 1)
ComponentProperty(12345, "\$Name", "Label1")
ComponentProperty(12345, "Text", "pet the kitty")
This sequence of events is then streamed to the user’s workspace via the RTC server.
III Results
III-A Coverage of App Inventor
Table I summarizes the current coverage Aptly provides for App Inventor functionality. We achieved our target of having one to one correspondence of the block language, with a few exceptions for new language features added by the App Inventor team since the Aptly project began.
TABLE I: Summary of Aptly’s block language coverage
Category Coverage %
Controls 100
Logic 100
Math 100
Text 94
Lists 100
Dictionaries 94
Colors 100
Procedures 75
Variables 80
Components 80
Total 96
TABLE II: Summary of Aptly Examples’ coverage of App Inventor functionality.
Feature Coverage %
Components 40.0
     Events 13.6
     Methods 6.4
     Properties 13.7
Coverage of App Inventor’s component library is more limited as shown in Table II. When building examples, we have aimed to cover the most common functionality observed in projects as studied by the App Inventor team [16]. Importantly, there is a very long tail of component usage that suggests that our initial set of example projects likely covers most use cases encountered by App Inventor users.
III-B Pilot User Study
To evaluate the effectiveness and user experience of our tool, Aptly, we conducted a study involving 10 high school students with the following demographics:
@@ III-B Pilot User Study
This diverse group was selected to yield insights into how students with different expertise levels and app creation familiarity engage with our tool. The study duration ranged from 50 to 80 minutes per participant. Initially, we introduced the students to Aptly, demonstrating how to create a basic app with interactive features such as a button and a label, and how to modify properties like color and font using the tool’s editing capabilities. We then assigned two distinct tasks to assess the tool’s utility and the participants’ ability to leverage it.
Figure 7: App Inventor Basic Tutorial: Digital Doodle
III-B1 Recreation Task
Students were shown “Digital Doodle” (Figure 7), a beginner-level app from App Inventor, and asked to recreate it using Aptly. This task was designed to observe whether participants could effectively translate a clear visual concept into a functional app through verbal descriptions, reflecting their understanding and communication of app design concepts. The task demonstrated the platform’s flexibility in understanding and executing a range of instructions, from detailed descriptions specifying colors and functionalities of buttons to succinct commands like “Draws with RGB.” This versatility highlights Aptly’s capacity to interpret and implement user input effectively, allowing users with no expert programming skills to successfully create functional apps.
Participants employed different strategies in app creation, some opting for a more incremental approach using Aptly’s editing features. For instance, one participant began with a basic layout and progressively added functionalities and aesthetic adjustments, such as alignment and color changes, through subsequent commands. This iterative process was particularly favored by participants with programming backgrounds, reflecting a workflow similar to traditional software development where features are built and refined over time. This suggests that Aptly can accommodate various user preferences and workflows, enhancing its usability across different skill levels and development styles. However, the study also uncovered challenges related to the precision of language used in commands. Some participants struggled to articulate their desired app functionalities clearly, leading to discrepancies between their intentions and the app’s behavior. For example, a participant’s command to “draw a line when I click a button” was too vague without specifying the color interaction, illustrating the current limitations of AI in interpreting user intent without explicit instruction.
III-B2 Scenario-Based Task

### v2

@@
Abstract.
This paper introduces Aptly, a platform designed to democratize mobile app development, particularly for young learners. Aptly integrates a Large Language Model (LLM) with App Inventor, enabling users to create apps using their natural language. User’s description is translated into a programming language that corresponds with App Inventor’s visual blocks. A preliminary study with high school students demonstrated the usability and potential of the platform. Prior programming experience influenced how users interact with Aptly. Participants identified areas for improvement and expressed a shift in perspective regarding programming accessibility and AI’s role in creative endeavors.
@@ Abstract.
Computational Action, Large Language Model, Block Programming, Mobile Application
journalyear: 2025copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706599.3720081isbn: 979-8-4007-1395-8/2025/04ccs: Human-centered computing User interface programming
Figure 1. The main Aptly interface is integrated with MIT App Inventor. On the top left there is “Project from description” button. Here the user typed in “Make an app that tracks what is in my fridge, I should able to add items and also delete them”.
1. Introduction
@@
Our core goal is to empower young people to develop technology worth making novel, digital solutions to problems young people face every day in their lives, their communities, and their world. Mobile applications serve as a powerful medium for such engagement given the widespread adoption of smartphones and tablets across diverse demographics, including under-represented communities (Napoli and Obar, 2014). However, the traditional app development process, which often requires a solid foundation in computer science, has historically excluded many aspiring creators from participating. Initiatives like App Inventor have democratized mobile app creation by introducing a user-friendly, visual programming environment. This block-based approach allows anyone, regardless of technical background, to develop applications by assembling geometric shapes representing code blocks (Wolber et al., 2015; Mohamad et al., 2011) The tool not only fosters creativity among novice programmers but also encourages a more inclusive approach to technological education and development (Bottino et al., 2020). Tissenbaum et al. call the process by which people leverage technology like App Inventor to make societal impact computational action (Tissenbaum et al., 2019).
@@
The advent of large language models (LLMs) (Vaswani et al., 2017) and their ability to generate code (Sarsa et al., 2022; Su et al., 2023) has opened new avenues for mobile app development. Capitalizing on this technological advancement, our research introduces Aptly, a platform that allows users to create mobile applications using natural language. For example, a user could simply state, “Make me an app that translates English to Spanish,” and Aptly would automatically generate a functional app based on this input. In a high level, the platform works as the following: The user describes their desired app or modification, the LLM translate that into a textual language that has a one-to-one corresponding with App Inventor’s blocks. Finally, the App Inventor platform compiles the blocks into a functional mobile app. This paper details the design of our platform, discussing the development of the Aptly Language and our method of integrating it with an LLM to produce the desired syntax. We conclude by presenting preliminary findings from its deployment among high school students.
2. User Interface
The Aptly user interface (Figure 1), integrated with App Inventor platform, introduces a new button labeled “Project from description”. When clicked, opens a window allowing the user to select a model for code generation and input a description of the app they wish to create. After detailing their app, users click “code it” to activate a large language model that generates the corresponding code. This process will be explored in greater detail later in the paper. Once the code is generated, it appears in a textbox on the right, where users can make manual adjustments as needed. After refining the code to their satisfaction, users press “make it” which automatically parses the code into a format compatible with App Inventor. The user is then redirected to the standard App Inventor interface, where they can continue developing the app using the designer and blocks features.
Users can describe the layout and specify the functionality of each user interface (UI) component within their mobile application. For example, users might design an application such as “Make an app with a text box and a button. When the button is clicked, the app says the words in the text box out loud. If the textbox is empty say “I have nothing to say””. Another instance is “Create an app with a button in the center and a label below the button that tracks the number of times the button is clicked.”. Additionally, Aptly supports the development of applications where the user does not need to specify detailed functionalities. For example, a user can simply request, “make a calculator app” and Aptly generates the application based on this high-level directive. This capability demonstrates Aptly’s ability to operate both as a detailed application development tool where specific UI behaviors are defined and as a more abstract tool that interprets and implements broad application concepts.
Users also have the capability to further refine their applications. The user can summon the pop-up window and specify the modifications they desire. For example, a user might request specific adjustments like “Make the font of the label three times larger” or more general changes such as “Make the button bigger” or even abstract alterations like “Change the color of the button to the color of a melon”. Aptly extends beyond mere automation by augmenting the creative process itself. As an intelligent agent powered by a large language model (LLM), Aptly assists users in achieving designs that surpass their initial conceptions. For example, the optimal color schemes for colorblind individuals might not be readily apparent to the average user. However, by using Aptly, a user can simply request that the UI color scheme be adjusted to be colorblind friendly, and the platform will automatically apply an appropriate palette based on its extensive knowledge base.
3. Platform Design
3.1. The Aptly Language
@@
Figure 2. An example of the Aptly code in Listing 1 and its correspondence to the App Inventor interface.
@@
Given these attributes, we hypothesize that generative models can be effectively tuned to interface with Aptly, leveraging its Python-like structure. However, it is important to note that, while Aptly draws inspiration from Python, it is distinct. Aptly has been specifically constrained to ensure a one-to-one correspondence with App Inventor: every valid App Inventor program is also a valid Aptly program and vice versa. See Listing 1 for an example Aptly program and Figure 2 to showcase how the code is represented as blocks.
Figure 3. When the user requests an application with its textual description, we automatically synthesize the prompt by adding several example pairs along with the desired application’s textual description. This constructed prompt is fed into one of OpenAI’s GPT-X models as an input, which outputs code that can be converted into a fully functional mobile application
3.2. Few-shot Prompt Engineering
LLMs operate based on “prompts” that are crucial inputs that significantly influence their performance. The effectiveness of these models hinges on the quality of the prompts, a practice known as prompt engineering (Reynolds and McDonell, 2021). A common method of prompt engineering involves “few-shot” prompts, where a model is given a limited number of solved task examples as part of its input. This approach is often interpreted as the model “learning” the task in real time from these few examples (Brown et al., 2020). Aptly uses Open AI’s LLMs, which are capable of generating computer code from natural language descriptions in several programming languages.
@@ 3.2. Few-shot Prompt Engineering
Aptly’s method for providing relevant prompts to accompany a description $D$ is to provide a set of example application descriptions paired with appropriate Aptly Code. We use semantic similarity to highlight the “appropriateness” of code and description. The method relies on having a way to automatically measure the similarity of text descriptions and the program. For computing the similarity of the text description and the program, we use embeddings of data elements, which are numerical representations of concepts converted to number sequences (Pennington et al., 2014). In our scenario, an embedding represents the semantic meaning of a natural language description or code. To measure the similarity of two items, one embeds them in the same vector space and takes the distance between the two embedding vectors as a measure of similarity (small distance implies high similarity, while large distance implies low similarity). We used the cosine distance, which reflects the angle between the vectors, to compute the similarity between two vectors.
@@
We employ OpenAI’s Text/Code Embedding model to generate embeddings from descriptions and associated code. These neural network models utilize Contrastive pre-training, a training method that functions by clustering in a vector space (Neelakantan et al., 2022). Specifically, it draws predefined positive examples (i.e., matching text and code) closer together, while pushing negative examples (i.e., contrasting text and code) apart. This technique enhances our model’s ability to identify and select the example pair most relevant to a user’s specified description. We feed the synthesized prompt into GPT, which outputs the Aptly Code corresponding to the user’s requested app description. We can then convert the generated code into App Inventor blocks to generate a fully functional application (Figure 3).
4. Pilot User Study
To evaluate the effectiveness and user experience of our tool, we conducted a study involving 10 high school students with the following demographics:
@@ 4. Pilot User Study
This group was selected to yield insights into how students with different expertise levels and app creation familiarity engage with our tool. The study duration ranged from 40 to 60 minutes per participant. Initially, we introduced the students to Aptly, demonstrating how to create a basic app with interactive features such as a button and a label, and how to modify properties like color and font using the tool’s editing capabilities. We then assigned two distinct tasks to assess the tool’s utility and the participants’ ability to leverage it.
Figure 4. App Inventor Basic Tutorial: Digital Doodle
4.1. Recreation Task
Students were shown “Digital Doodle” (Figure 4), a beginner-level app from App Inventor, and asked to recreate it using Aptly. This task was designed to observe whether participants could effectively translate a clear visual concept into a functional app through verbal descriptions, reflecting their understanding and communication of app design concepts. The task required the platform’s flexibility in understanding and executing a range of instructions, from detailed descriptions specifying colors and functionalities of buttons (e.g., An app with 3 buttons (one red, one blue, one green), and a pen drawing feature with a blank “drawing” space beneath it. When the mouse is dragged in the blank space, create a line where the drag is. When red is clicked, the pen color is red. When blue is clicked, the pen color is blue. When green is clicked, the pen color is green.) to succinct commands like “Draws with RGB.”.
Participants employed different strategies in app creation, some opting for a more incremental approach using Aptly’s editing features. For instance, one participant began with a basic layout and progressively added functionalities and aesthetic adjustments, such as alignment and color changes, through subsequent commands. This iterative process was particularly favored by participants with programming backgrounds, reflecting a workflow similar to traditional software development where features are built and refined over time. The study also uncovered challenges related to the precision of language used in commands. Some participants struggled to articulate their desired app functionalities clearly, leading to discrepancies between their intentions and the app’s behavior. For example, a participant’s command to “draw a line when I click a button” was too vague without specifying the color interaction.
4.2. Scenario-Based Task